OpenStudy (anonymous):

A spinner with three congruent sections is spun with the results as shown. What is P(1)? 1 3 3 2 1 2 1 1 3 3 2 1 1 1 2 2 2 3 3 1 3

OpenStudy (anonymous):

The number of times that a spin comes up 1 divided by the total number of spins. P(1) = 8/21

OpenStudy (anonymous):

can you help me with one more question please Two urns each contain green balls and blue balls. Urn I contains 4 green balls and 6 blue balls, and Urn II contains 6 green balls and 2 blue balls. A ball is drawn at random from each urn. What is the probability that both balls are blue?

OpenStudy (anonymous):

its rather 2/51 3/20 or 4/153

OpenStudy (anonymous):

and i got 3/20

OpenStudy (cherylim23):

Pr(blue in urn 1) = 6/(6+4) = 6/10 = 3/5 Pr(blue in urn 2) = 2(6+2) = 2/8 = 1/4 Multiply these two probabilities together, as you want them to occur together, 3/5*1/4 = 3/20 so yes you're right

OpenStudy (anonymous):

you can think of it as two tables with different colored squares. 10 balls total in urn 1. P1(blue)= 6/10 8 balls total in urn 2. P2(blue)=2/8 Multiply together 6/10 * 2/8=0.15